An SRS of 55 lawyers from Reno, Nevada had a mean hourly salary of x= $310 and standard deviation s- $75 a. construct a 95% confidence interval for the mean hourly salary of all Reno lawyers b.find the margin of error in part a c. how large a sample would be necessary to have a margin of error of $25 with a 95% confidence level?
Answer #1

Solution:

n = 55

\bar{x} = 310

s = 75

( a )

z = 1.96

95% confidence interval = 310 \pm 1.96 (75/\sqrt{55})

= 310 \pm 19.8214

= ( 290.1786 , 329.8214 )

( b )

margin of error =  1.96 (75/\sqrt{55})

= 19.8214

( c )

ME = 25

we know that

ME = z ( s / \sqrt{n} )

25 = 1.96 ( 75 / \sqrt{n} )

\sqrt{n} = 1.96*75/25

\sqrt{n} = 5.88

n = 34.57

n \approx 35